OpenStudy (anonymous):

What integral represents the volume of the solid formed by revolving the region bounded by the graphs of y = x^3 y = 1 and x = 2 about the line x = 2

OpenStudy (mle):

Hello @Yaros and Welcome to OpenStudy! Thanks for asking a Qualified Help Question first try graphing out what you have been given so that you have a visual representation of what you are looking for :)

OpenStudy (michele_laino):

your problem can be represented by this drawing: 

OpenStudy (michele_laino):

now I make this traslation: \[\Large \left\{ \begin{gathered} x = X + 2 \hfill \\ y = Y \hfill \\ \end{gathered} \right.\] where X, Y are the new coordinates, furthermore, we can see that the origin of the plane X-Y is lòocated at point x=2,y=0

OpenStudy (michele_laino):